# Probing the equilibration of the QCD matter created in heavy-ion collisions with dileptons

### v1

@@ III.2.1 M-dependent observables
The research in this paper focuses mainly on thermal dileptons. Therefore, for the calculations of dilepton production and dilepton flow observables, only cells where the local temperature exceeds the chemical freeze-out temperature ( GeV) Adamczyk et al. (2017) are considered.
@@
In Fig. 3, the dilepton yield is presented as a function of invariant mass in the 0-5% centrality class at TeV Pb+Pb collisions at the LHC. The dilepton production contains different contributions including the pre-equilibrium stage – which is analyzed under three kinds of chemical equilibrium scenarios – the hydrodynamics stage (labeled as thermal), and the Drell-Yan process Drell and Yan (1970). At LHC collision energies, the Drell-Yan process is one of the main sources of dilepton production at high invariant masses. It is an important test of the Standard Model, and has been measured for high mass dileptons in p+p collisions at the LHC Aad et al. (2014). However, in the low and intermediate invariant mass regions in heavy-ion collisions, both measurements and theoretical calculations of the Drell-Yan process are scarce.
@@
In this work we utilize the DYTurbo package Camarda et al. (2020) to estimate the production of dileptons from the Drell-Yan process at low and intermediate invariant masses. The DYTurbo program is designed to calculate the Drell-Yan process at NLO in p+p collisions, it requires adjustments for A+A collisions. These adjustments include considering the nuclear parton distribution functions nPDFs, specifically EPPS16nlo-CT14nlo-Pb208 Eskola et al. (2017), as well as accounting for the number of binary collisions as determined by the MCGlauber modelLoizides et al. (2018), and the inelastic cross-section from p+p collisions, in the cross-section calculations performed by DYTurbo:
@@
In Fig. 3, the band obtained with DYTurbo representing dilepton production from Drell-Yan arises from the variation in the factorization and renormalization scales. The upper limit corresponds to the factorization and renormalization scales being twice the invariant mass of the lepton pair, while the lower limit corresponds to half the mass Coquet et al. (2021). In perturbative QCD calculations, parton distribution functions and fragmentation functions are typically limited from below by a practical scale of the order of GeV. This is also the case for DYTurbo and this explains the absence of results at lower invariant masses. Note that in addition to using DYTurbo to estimate dilepton production from the Drell-Yan process, the experimental community also often relies on Pythia6 Sjostrand et al. (2006); Adamczyk et al. (2015a, b). Results obtained with DYTurbo and Pythia6 will be compared in Appendix A.
@@
In Fig. 3, it is observed that thermal dilepton production dominates in the low invariant mass range compared to other dilepton sources. This is because dileptons at low invariant mass are primarily produced in relatively low-temperature regions, which correspond to the later stages of heavy-ion collisions. As the invariant mass increases, the contribution of dileptons from the pre-equilibrium stage gradually increases and becomes the dominant source over thermal dileptons for GeV. Keep in mind that dilepton production from the Drell-Yan process in the invariant mass region from 1 to 4 GeV exhibits large theoretical systematic uncertainty, even though the Drell-Yan process can be accurately measured at high invariant mass ( GeV)Aad et al. (2014). It is also found that dilepton production through Drell-Yan is consistently smaller than the pre-equilibrium contribution until GeV. Taken together, those observations indicate that dilepton production from the pre-equilibrium stage is observable in the intermediate invariant mass region, and that any quantitative differences with previous results Kasmaei and Strickland (2019); Churchill et al. (2021); Coquet et al. (2021); Garcia-Montero et al. (2024b) highlight the importance of realistic initial states, of complete emission rates, and of complete modeling. Regarding the effect of the chemical equilibrium, it is observed that dileptons induced from the pre-equilibrium phase with a higher-power of the SF chemical equilibrium factor will be suppressed. In addition, the higher mass regions directly correlates with high temperatures Churchill et al. (2024a). Thus, dileptons production in the intermediate invariant mass region will be sensitive to the early chemical equilibrium, as shown on Fig. 3.
@@
By summing over the contributions from the hydrodynamics and pre-equilibrium stages first shown in Fig. 3, we plot the total thermal dilepton production as a function of invariant mass in the 0-20%, 20-40%, and 40-60% centrality classes at LHC collision energy in panels (a), (c), and (e) of Fig. 4. Each panel includes three different curves corresponding to three scenarios of chemical equilibrium. Panels (b), (d), and (f) present the ratios between the linear and quadratic orders to the instant chemical equilibrium. In each centrality class, the effect of chemical equilibrium primarily influences the intermediate invariant mass region. From the panels of ratios, it is clearer to see that higher orders of chemical equilibrium suppress dilepton production more significantly, especially at an invariant mass of 4-5 GeV, where only 40% of dilepton production with quadratic order suppression factor is observed compared to the instant chemical equilibrium. Additionally, the effect of chemical equilibrium generally increases with the centrality class and the difference between instant and quadratic chemical equilibrium becomes more distinct in peripheral collisions. This can be attributed to the relative contribution of the pre-equilibrium stage which grows in peripheral collisions.
@@
We now discuss the dilepton anisotropic flow coefficients. In all of the flow calculations in this work, we do not include the contribution of Drell-Yan dileptons. Their elliptic flow can be assumed to vanish, and their yield can however be calculated, either with DYTurbo or Pythia (see the Appendix) and subtracted from measured signals, as done for the pQCD production of real photons Esha (2024). Clearly, a direct measurement in p+p collisions would be advantageous, as it would largely alleviate the scale uncertainty found in DY calculations. Nevertheless, we provide the following estimates: for the invariant mass-dependent flow (Fig. 5) at M = 2 GeV, including the DY contribution reduces the elliptic flow by an amount varying between 2% and 8% (depending on the choice of scales) in the 0-20% centrality bin. In the 20-40% and 40-60% bins, the reduction ranges from 2% to 11% and from 3% to 16%, respectively.
@@
Fig. 5 and 6 show the -integrated thermal and total dilepton elliptic flow and triangular flow , in the 0-20%, 20-40%, and 40-60% centrality classes. The different order of chemical equilibrium (power of the SF) is shown in the total dilepton flow coefficient. In a given centrality class, the elliptic flow coefficient initially increases, reaches a peak at about GeV, then starts to decrease as the invariant mass increases, owing to the smaller momentum anisotropy of the collision system at the early stage. After considering the contribution of the pre-equilibrium stage, the total dilepton will be suppressed due to the almost zero momentum anisotropy in pre-equilibrium stage. Particularly, the total dilepton flow at the invariant mass larger than 2 GeV region, where the pre-equilibrium is dominant, shows a more significant suppression. When focusing on the effect of chemical equilibrium, it displays a trend opposite to that seen in the dilepton yields. For the higher order of the suppression factor, the total dilepton flow is closer to that of thermal dileptons. This is because a higher-order suppression factor will further suppress the contribution of dilepton production at the early stage, so the relative contribution of later proper times in pre-equilibrium will be enhanced as the flow builds up as time grows.
@@
In addition to studying the invariant mass dependence of dilepton production and of flow, it is equally revealing to examine their transverse momentum () dependence because different invariant mass windows can provide insight into various periods of the QGP evolution history. Fig. 7 displays the dilepton production as a function of transverse momentum in different invariant mass windows: (a) GeV and (b) GeV, within the 0-5% centrality class in TeV Pb+Pb collisions. Within the invariant mass window of GeV, which predominantly originates from the low-temperature region of the QGP medium, the dilepton spectrum is consequently dominated by the thermal production. However, beginning in the intermediate region, dileptons from the pre-equilibrium stage become dominant. Additionally, it is observed that the slope of dilepton production in pre-equilibrium is flatter, reflecting the natural property for dileptons in pre-equilibrium to possess higher transverse momentum. In the GeV invariant mass window, the pre-equilibrium dilepton production is seen to be comparable to the thermal dilepton production in the low region. Especially in the case of instant chemical equilibrium, the pre-equilibrium dilepton production consistently exceeds thermal dilepton production and Drell-Yan dileptons. This indicates that by carefully selecting the invariant mass window it is possible to probe the properties of the pre-equilibrium stage using dilepton production. Furthermore, in this centrality class the transverse momentum dependence of dilepton production is only moderately influenced by the effects of chemical equilibrium.
@@
In Fig. 8, the dilepton flow is presented as a function of transverse momentum () within the invariant mass windows of GeV and GeV, for the 0-20% and 20-40% centrality classes. The total dilepton flow is also calculated assuming different values of the suppression factor, as explained previously. In the low invariant mass range, shown in panels (a) and (c), which predominantly originates from the low-temperature region of the QGP medium, thermal dileptons produce a sizable elliptic flow. However, after accounting for pre-equilibrium dileptons, the total dilepton flow is significantly suppressed relative to the thermal dilepton flow, particularly at higher transverse momenta , where pre-equilibrium dileptons are known to dominate, as demonstrated in Fig. 7. In the intermediate invariant mass range (panels (b) and (d)), it is even more obvious that the total dilepton elliptic flow experiences a noticeable reduction due to the dominance of pre-equilibrium dileptons in this invariant mass range. Furthermore, it is interesting to note that the dilepton total elliptic flow is very sensitive to the effects of chemical equilibrium. The ratio of total dilepton flow with a linear and quadratic SF to values assuming instant chemical equilibrium are approximately 1.5 and 2.5, respectively, in the 20-40% centrality class of panel (c). This indicates that, considering all centrality classes, dilepton flow do provide a good potential for quantifying the effects of chemical equilibrium in the pre-equilibrium stage of heavy ion collisions. The same conclusion can also be obtained from the analysis of dilepton triangular flow in Fig. 9.
@@ IV Conclusion
In this work, an analysis of thermal dilepton production and anisotropic flow is performed in Pb+Pb collisions at a LHC collision energy of TeV. A state-of-the-art multistage model, comprising the QCD-based IP-Glasma, the out-of-equilibrium KøMPøST , the relativistic hydrodynamic MUSIC, and the hadronic afterburner UrQMD, is used to simulate the entire dynamic evolution of relativistic heavy ion collisions.After calibrating the model, we investigate dilepton production at the different stages, using dilepton production rates complete at NLO in the strong coupling. It is found that thermal dileptons are the dominant source at low invariant mass, and pre-equilibrium dileptons overtake thermal dileptons to become the main contribution at intermediate invariant mass ranges, exceeding the dilepton production from the Drell-Yan process. The effect of chemical equilibrium have been further studied. We found that the departure from chemical equilibrium can suppress dilepton production and enhance dilepton anisotropic flow.
This paper also studies the transverse momentum dependence of dilepton observables. One interesting result is that the transverse momentum spectrum of pre-equilibrium dileptons can dominate over that of other sources, in the intermediate invariant mass window. This suggests that dilepton production can provide a window to explore the properties of the pre-equilibrium stage after selecting the appropriate invariant mass range. In the low invariant mass range, the flow results at intermediate centrality shown in Fig. 8 and 9, show a sensitivity to the chemical equilibrium effect. However, a proper assessment in this mass range will require the inclusion of dileptons from other hadronic sources. Indeed, the dileptons from reactions involving composite hadrons are also a very important source for low invariant masses dileptons Rapp and Wambach (2000). Finally, adapting the dilepton rates to account for the various corrections discussed earlier in the text will be important for high precision future studies. Incorporating those many aspects in a comprehensive and detailed dilepton analysis can help us gain a deeper understanding of the dynamics of heavy ion collisions and on the nature of the QCD phase diagram.
@@ Acknowledgements.
It is a pleasure to thank Greg Jackson for contributions that led to this work, and for valuable discussions. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada, and in part by US National Science Foundation (NSF) under grant number OAC-2004571. Computations were made on the Béluga super-computer system from McGill University, managed by Calcul Québec and by the Digital Research Alliance of Canada.Appendix A Drell-Yan dileptons from DYTurbo and Pythia6
We compare the predictions of DYTurbo against those from Pythia6, for the production of dileptons via the Drell-Yan process. In Fig. 10, the dilepton production from the Drell-Yan process estimated from the DYTurbo model and Pythia6 event generator in 0-5% centrality Pb+Pb collisions at 5.02 TeV is shown. DYTurbo has the same set up as that used in Fig. 7. In the Pythia6 calculation, the parameter for the width of the Gaussian primordial distribution is taken as 0.95 (PARP(9) = 0.95), the square of the momentum transfer equals the square of the dilepton pair mass (MSTP(32) = 4), and it uses the default PDF sets CTEQ5L. The bands in Pythia6 result from varying the K factor, which multiplies the differential cross section for hard parton–parton processes to effectively account for higher order corrections, from 1.0 to 1.8 (PARP(31) = 1.0 - 1.8). It is observed that dilepton production from the Drell-Yan process exhibits significant model uncertainty in both cases. Pythia6 predicts a higher dilepton yield compared to the DYTurbo calculation when the invariant mass () GeV. But there is a clear region of overlap between both approach to the calculation of DY dileptons. Accurate estimates of the Drell-Yan process will significantly affect the extraction of dilepton production from the pre-equilibrium stage, and this highlights the importance of measuring the Drell-Yan process at intermediate invariant masses.

### v2

@@ III.2.1 M-dependent observables
The research in this paper focuses mainly on thermal dileptons. Therefore, for the calculations of dilepton production and dilepton flow observables, only cells where the local temperature exceeds the chemical freeze-out temperature [ GeV] Adamczyk et al. (2017) are considered.
@@
In Fig. 3, the dilepton yield is presented as a function of invariant mass in the 0-5% centrality class at TeV Pb+Pb collisions at the LHC. The dilepton production contains different contributions including the preequilibrium stage – which is analyzed under three kinds of chemical equilibrium scenarios – the hydrodynamics stage (labeled as thermal), and the Drell-Yan process Drell and Yan (1970). At LHC collision energies, the Drell-Yan process is one of the main sources of dilepton production at high invariant masses. It is an important test of the standard model, and has been measured for high mass dileptons in p+p collisions at the LHC Aad et al. (2014). However, in the low and intermediate invariant mass regions in heavy-ion collisions, both measurements and theoretical calculations of the Drell-Yan process are scarce.
@@
In this work we utilize the DYTurbo package Camarda et al. (2020) to estimate the production of dileptons from the Drell-Yan process at low and intermediate invariant masses within a rapidity window . The DYTurbo program is designed to calculate the Drell-Yan process at NLO in p+p collisions, so it requires adjustments for A+A collisions. These adjustments include considering the nuclear parton distribution functions nPDFs (EPPS16nlo-CT14nlo-Pb208 Eskola et al. (2017)), as well as accounting for the number of binary collisions as determined by the Monte Carlo Glauber modelLoizides et al. (2018), and the inelastic cross section from p+p collisions, in the cross-section calculations performed by DYTurbo:
@@
In Fig. 3, the band obtained with DYTurbo representing dilepton production from Drell-Yan arises from the variation in the factorization and renormalization scales. The upper limit corresponds to the factorization and renormalization scales being twice the invariant mass of the lepton pair, while the lower limit corresponds to half the mass Coquet et al. (2021). In perturbative QCD calculations, parton distribution functions and fragmentation functions are typically limited from below by a practical scale of the order of GeV. This is also the case for DYTurbo and this explains the absence of results at lower invariant masses. Note that in addition to using DYTurbo to estimate dilepton production from the Drell-Yan process, the experimental community also often relies on Pythia6 Sjostrand et al. (2006); Adamczyk et al. (2015a, b). Results obtained with DYTurbo and Pythia6 will be compared in the Appendix A.
@@
In Fig. 3, it is observed that thermal dilepton production dominates in the low invariant mass range compared to other dilepton sources. This is because dileptons at low invariant mass are primarily produced in relatively low-temperature regions, which correspond to the later stages of heavy-ion collisions. As the invariant mass increases, the contribution of dileptons from the preequilibrium stage gradually increases and becomes the dominant source over thermal dileptons for GeV. Keep in mind that dilepton production from the Drell-Yan process in the invariant mass region from 1 to 4 GeV exhibits large theoretical systematic uncertainty, even though the Drell-Yan process can be accurately measured at high invariant mass ( GeV) Aad et al. (2014). It is also found that dilepton production through Drell-Yan is consistently smaller than the preequilibrium contribution until GeV. Taken together, those observations indicate that dilepton production from the preequilibrium stage is observable in the intermediate invariant mass region, and that any quantitative differences with previous results Kasmaei and Strickland (2019); Churchill et al. (2021); Coquet et al. (2021); Garcia-Montero et al. (2024b) highlight the importance of realistic initial states, of complete emission rates, and of complete modeling. Regarding the effect of the chemical equilibrium, it is observed that dileptons induced from the preequilibrium phase with a higher-power of the SF chemical equilibrium factor will be suppressed. In addition, the higher mass regions directly correlates with high temperatures Churchill et al. (2024a). Thus, dilepton production in the intermediate invariant mass region will be sensitive to the early chemical equilibrium, as shown in Fig. 3.
@@
By summing over the contributions from the hydrodynamics and preequilibrium stages first shown in Fig. 3, we plot the total thermal dilepton production as a function of invariant mass in the 0-20%, 20-40%, and 40-60% centrality classes at LHC collision energy in panels (a), (c), and (e) of Fig. 4. Each panel includes three different curves corresponding to three scenarios of chemical equilibrium. Panels (b), (d), and (f) present the ratios between the linear and quadratic orders to the instant chemical equilibrium. In each centrality class, the effect of chemical equilibrium primarily influences the intermediate invariant mass region. From the panels of ratios, it is clearer to see that higher orders of chemical equilibrium suppress dilepton production more significantly, especially at an invariant mass of 4-5 GeV, where only 40% of dilepton production with quadratic order suppression factor is observed compared to the instant chemical equilibrium. Additionally, the effect of chemical equilibrium generally increases with the centrality class and the difference between instant and quadratic chemical equilibrium becomes more distinct in peripheral collisions. This can be attributed to the relative contribution of the preequilibrium stage which grows in peripheral collisions.
@@
We now discuss the dilepton anisotropic flow coefficients. In all of the flow calculations in this work, we do not include the contribution of Drell-Yan dileptons. Their elliptic flow can be assumed to vanish, and their yield can, however, be calculated, either with DYTurbo or Pythia (see the Appendix) and subtracted from measured signals, as done for the pQCD production of real photons Esha (2024). Clearly, a direct measurement in p+p collisions would be advantageous, as it would largely alleviate the scale uncertainty found in DY calculations. Nevertheless, we provide the following estimates: for the invariant mass-dependent flow (Fig. 5) at M = 2 GeV, including the DY contribution reduces the elliptic flow by an amount varying between 2% and 8% (depending on the choice of scales) in the 0-20% centrality bin. In the 20-40% and 40-60% bins, the reduction ranges from 2% to 11% and from 3% to 16%, respectively.
@@
Figures. 5 and 6 show the -integrated thermal and total dilepton elliptic flow and triangular flow , in the 0-20%, 20-40%, and 40-60% centrality classes. The different order of chemical equilibrium (power of the SF) is shown in the total dilepton flow coefficient. In a given centrality class, the elliptic flow coefficient initially increases, reaches a peak at about GeV, then starts to decrease as the invariant mass increases, owing to the smaller momentum anisotropy of the collision system at the early stage. After considering the contribution of the preequilibrium stage, the total dilepton will be suppressed due to the almost zero momentum anisotropy in the preequilibrium stage. Particularly, the total dilepton flow at the invariant mass larger than 2 GeV region, where the preequilibrium is dominant, shows a more significant suppression. When focusing on the effect of chemical equilibrium, it displays a trend opposite to that seen in the dilepton yields. For the higher order of the suppression factor, the total dilepton flow is closer to that of thermal dileptons. This is because a higher-order suppression factor will further suppress the contribution of dilepton production at the early stage, so the relative contribution of later proper times in preequilibrium will be enhanced as the flow builds up as time grows.
@@
In addition to studying the invariant mass dependence of dilepton production and of flow, it is equally revealing to examine their transverse momentum () dependence because different invariant mass windows can provide insight into various periods of the QGP evolution history. Figure. 7 displays the dilepton production as a function of transverse momentum in different invariant mass windows: (a) GeV and (b) GeV, within the 0-5% centrality class in TeV Pb+Pb collisions. Within the invariant mass window of GeV, which predominantly originates from the low-temperature region of the QGP medium, the dilepton spectrum is consequently dominated by the thermal production. However, beginning in the intermediate region, dileptons from the preequilibrium stage become dominant. Additionally, it is observed that the slope of dilepton production in preequilibrium is flatter, reflecting the natural property for dileptons in preequilibrium to possess higher transverse momentum. In the GeV invariant mass window, the preequilibrium dilepton production is seen to be comparable to the thermal dilepton production in the low region. Especially in the case of instant chemical equilibrium, the preequilibrium dilepton production consistently exceeds thermal dilepton production and Drell-Yan dileptons. This indicates that by carefully selecting the invariant mass window it is possible to probe the properties of the preequilibrium stage using dilepton production. Furthermore, in this centrality class the transverse momentum dependence of dilepton production is only moderately influenced by the effects of chemical equilibrium.
@@
In Fig. 8, the dilepton flow is presented as a function of transverse momentum () within the invariant mass windows of GeV and GeV, for the 0-20% and 20-40% centrality classes. The total dilepton flow is also calculated assuming different values of the suppression factor, as explained previously. In the low invariant mass range, shown in panels (a) and (c), which predominantly originates from the low-temperature region of the QGP medium, thermal dileptons produce a sizable elliptic flow. However, after accounting for preequilibrium dileptons, the total dilepton flow is significantly suppressed relative to the thermal dilepton flow, particularly at higher transverse momenta , where preequilibrium dileptons are known to dominate, as demonstrated in Fig. 7. In the intermediate invariant mass range [panels (b) and (d)], it is even more obvious that the total dilepton elliptic flow experiences a noticeable reduction due to the dominance of preequilibrium dileptons in this invariant mass range. Furthermore, it is interesting to note that the dilepton total elliptic flow is very sensitive to the effects of chemical equilibrium. The ratios of total dilepton flow with a linear and quadratic SF to values assuming instant chemical equilibrium are approximately 1.5 and 2.5, respectively, in the 20-40% centrality class of panel (c). This indicates that, considering all centrality classes, dilepton flow does provide a good potential for quantifying the effects of chemical equilibrium in the preequilibrium stage of heavy ion collisions. The same conclusion can also be obtained from the analysis of dilepton triangular flow in Fig. 9.
@@ IV Conclusion
In this work, an analysis of thermal dilepton production and anisotropic flow is performed in Pb+Pb collisions at a LHC collision energy of TeV. A state-of-the-art multistage model, comprising the QCD-based IP-Glasma, the out-of-equilibrium KøMPøST , the relativistic hydrodynamic MUSIC, and the hadronic afterburner UrQMD, is used to simulate the entire dynamic evolution of relativistic heavy ion collisions. After calibrating the model, we investigate dilepton production at the different stages, using dilepton production rates complete at NLO in the strong coupling. It is found that thermal dileptons are the dominant source at low invariant mass, and preequilibrium dileptons overtake thermal dileptons to become the main contribution at intermediate invariant mass ranges, exceeding the dilepton production from the Drell-Yan process. KøMPøST was used to model the approach to kinetic equilibrium, and the effect of chemical equilibrium has been incorporated therein. We find that the departure from chemical equilibrium can suppress dilepton production and enhance dilepton anisotropic flow.
This paper also studies the transverse momentum dependence of dilepton observables. One interesting result is that the transverse momentum spectrum of preequilibrium dileptons can dominate over that of other sources, in the intermediate invariant mass window. This suggests that dilepton production can provide a window to explore the properties of the preequilibrium stage after selecting the appropriate invariant mass range. In the low invariant mass range, the flow results at intermediate centrality shown in Figs. 8 and 9, show a sensitivity to the chemical equilibrium effect. However, a proper assessment in this mass range will require the inclusion of dileptons from other hadronic sources. Indeed, the dileptons from reactions involving composite hadrons are also a very important source for low invariant masses dileptons Rapp and Wambach (2000). Finally, adapting the dilepton rates to account for the various corrections discussed earlier in the text will be important for high precision future studies. Incorporating those many aspects in a comprehensive and detailed dilepton analysis can help us gain a deeper understanding of the dynamics of heavy ion collisions and the nature of the QCD phase diagram.
@@ Acknowledgements.
It is a pleasure to thank Greg Jackson for contributions that led to this work, for valuable discussions, and for a thorough reading of an earlier draft of this paper. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada, and in part by US National Science Foundation (NSF) under grant number OAC-2004571. Computations were made on the Béluga super-computer system from McGill University, managed by Calcul Québec and by the Digital Research Alliance of Canada.Appendix A Drell-Yan dileptons from DYTurbo and Pythia6
We compare the predictions of DYTurbo against those from Pythia6, for the production of dileptons via the Drell-Yan process. In Fig. 10, the dilepton production from the Drell-Yan process estimated from the DYTurbo model and Pythia6 event generator in 0-5% centrality Pb+Pb collisions at 5.02 TeV is shown. DYTurbo has the same set up as that used in Fig. 7. In the Pythia6 calculation, the parameter for the width of the Gaussian primordial distribution is taken as 0.95 [PARP(9) = 0.95], the square of the momentum transfer equals the square of the dilepton pair mass [MSTP(32) = 4], and it uses the default PDF sets CTEQ5L. The bands in Pythia6 result from varying the K factor, which multiplies the differential cross section for hard parton–parton processes to effectively account for higher order corrections, from 1.0 to 1.8 [PARP(31) = 1.0 - 1.8]. It is observed that dilepton production from the Drell-Yan process exhibits significant model uncertainty in both cases. Pythia6 predicts a higher dilepton yield compared to the DYTurbo calculation when the invariant mass GeV. But there is a clear region of overlap between both approach to the calculation of DY dileptons. Accurate estimates of the Drell-Yan process will significantly affect the extraction of dilepton production from the preequilibrium stage, and this highlights the importance of measuring the Drell-Yan process at intermediate invariant masses.